# Constraining Participation: Affordances of Feedback Features in Interfaces to Large Language Models

### v1

@@ Abstract.
Large language models (LLMs) are now accessible to anyone with a computer, a web browser, and an internet connection via browser-based interfaces, shifting the dynamics of participation in AI development. This paper examines the affordances of interactive feedback features in ChatGPT’s interface, analysing how they shape user input and participation in LLM iteration. Drawing on a survey of ChatGPT users and applying the mechanisms and conditions framework of affordances, we demonstrate that these features encourage simple, frequent, and performance-focused feedback while discouraging collective input and discussions among users. We argue that this feedback format significantly constrains user participation, reinforcing power imbalances between users, the public, and companies developing LLMs. Our analysis contributes to the growing body of literature on participatory AI by critically examining the limitations of existing feedback processes and proposing directions for their redesign. To enable more meaningful public participation in AI development, we advocate for a shift away from processes focused on aligning model outputs with specific user preferences. Instead, we emphasise the need for processes that facilitate dialogue between companies and diverse ‘publics’ about the purpose and applications of LLMs. This approach requires attention to the ongoing work of infrastructuring—creating and sustaining the social, technical, and institutional structures necessary to address matters of concern to groups impacted by AI development and deployment.
@@
copyright: none
@@ 1. Introduction
The deployment of ChatGPT by OpenAI in 2022 marked a significant shift in how large language models (LLMs) are made available to the public. ChatGPT’s browser-based interface allows broad access to interactions with LLMs through text prompts, expanding user participation in the iterative development of LLMs. Previously, companies such as OpenAI and Google took a more cautious approach. OpenAI used application programming interfaces (APIs) and an iterative approach to deploy the GPT-2 and GPT-3 family of models (Solaiman et al., 2019; Brundage et al., 2022), limiting access to developers and researchers. Similarly, Google announced versions of the Language Model for Dialogue Applications (LaMDA) in 2021 and 2022 but did not make it publicly available through an interface or model release (Thoppilan et al., 2022). Following ChatGPT’s deployment, numerous other LLMs were made accessible via browser-based interfaces (e.g., Bard/Gemini and Claude.ai) or open releases (e.g., LLaMA).
@@
    What are the affordances of interactive feedback features in the ChatGPT interface?
@@
    How do those affordances shape the nature and scope of user input and participation in the iteration of LLMs?
To address these questions, we first surveyed ChatGPT users who have submitted feedback. We then drew on the survey insights to analyse interactive features in the ChatGPT interface that allow users to submit feedback, using Davis’s (Davis, 2020) mechanisms and conditions framework of affordances. Our analysis indicates that the ChatGPT interface encourages simple, frequent, individualised, and unidirectional feedback on performance-related concerns while discouraging feedback from collectives or discussions among users.
We discuss implications for (re)designing feedback processes for LLMs. First, we explore deliberative approaches that might enable bidirectional communication, changes in perspectives over time, and discussion about social or ethical issues among groups of users. However, we recognise that deliberation on rulesets for model outputs will not fundamentally alter the power dynamics between users, the public, and platforms. Second, we consider the “infrastructuring” (Le Dantec and DiSalvo, 2013) required to support and sustain more expansive feedback channels between the public and technology companies. We conclude by urging companies deploying LLMs to shift their focus from aligning model outputs with user preferences to fostering sustained dialogue with diverse ‘publics’ about the purposes and applications of these technologies. This approach requires careful attention to the ongoing work of infrastructuring participation—creating and maintaining the social, technical, and institutional structures necessary for genuine public engagement in AI development. We must move beyond preference elicitation before LLMs become ubiquitous infrastructure—systems that are embedded in other structures, transparent in use, and only visible upon breakdown (Star and Ruhleder, 1996).
@@ 2.1. Alignment and Participatory Approaches
Interest in participatory approaches for aligning LLMs is part of a broader participatory turn in AI research, design, and development (Delgado et al., 2023). Participatory Al projects involve end users and those impacted by AI systems in the design and development of those systems (Cooper and Zafiroglu, 2024), drawing on long-standing traditions in other fields, such as participatory design (Muller and Kuhn, 1993; Sundblad, 2011) and participatory planning (Arnstein, 1969; Kelty, 2020). The turn to participation in AI research, design and development has included papers, panels and workshops on participatory machine learning (ML) at prominent computer science conferences (Kulynych et al., 2020; Zytko et al., 2022) and calls for public participation in AI in recent policy frameworks (Farthing et al., 2021; U.S. Office of Science and Technology Policy, 2022; Tabassi, 2023).
Participatory approaches promise more “democratic” AI systems (Kulynych et al., 2020) and more “pluralistic” alignment (Kirk et al., 2024). However, there is a broad spectrum of participatory practices in AI design, from informing or consulting participants to equitable partnerships with participants (Birhane et al., 2022; Corbett et al., 2023; Delgado et al., 2023). This spectrum of participatory practices also existed in participatory design, participatory planning, and other antecedent fields. For the development of large-scale AI systems, participation has often taken passive and involuntary forms through the representation of groups of individuals in datasets used to train models (Denton et al., 2020; Sloane et al., 2022). While feedback features in interfaces to LLMs provide a more active and voluntary form of user participation, their emergence as the primary mode for the public to influence development processes has raised questions about the conceptual leap from structured preference elicitation to genuine participatory engagement (Feffer et al., 2023b).
Additionally, there is a significant gap between the scale of participatory methods and the training and deployment of recent AI systems (Diaz and Madaio, 2023). Surveys of participatory approaches in computing research (Cooper et al., 2022) and Participatory AI projects (Delgado et al., 2023) indicate that engagements have primarily been localised, focusing on one or a few groups or communities with shared characteristics. In contrast, recent deployments of LLMs through browser-based interfaces are accessible to anyone with access to a computer, a web browser, and an internet connection.
@@ 2.2. Interactive Feedback Features in Interfaces to LLMs
Feedback has a rich history rooted in cybernetics, which introduced the idea of feedback loops, where the output of a system is fed back as input, and queried the implications of feedback loops for modelling dynamical systems over time (Wiener, 1948). The concept of feedback loops influenced subsequent AI research directions (Minsky, 1961) and remains fundamental to AI systems that learn over time. Concurrently, ‘feedback’ has also become a general term referring to information provided by an observer (e.g., users or other stakeholders) assessing a system’s performance.
Reinforcement Learning from Human Feedback (RLHF) demonstrates how the two types of feedback are combined in modern AI development. RLHF begins with evaluative feedback, when a model generates two responses to a given prompt and human raters determine which response is better than the other according to criteria. These human assessments are then used to create a cybernetic feedback loop: an RL algorithm uses the assessments to predict how a human would rate other responses, and an LLM is fine-tuned on the resulting reward model (Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022). This approach has become central to aligning pre-trained LLMs with specific preferences (Lambert et al., 2023). However, integrating human feedback into the learning process increases the time to train a model and the labour costs involved (Maslej et al., 2024). RLHF requires large, diverse groups of human raters to assess prompt completions before fine-tuning the model in the final stage (Prabhakaran et al., 2021).
The importance of RLHF for aligning LLMs has increased the demand for human feedback data (Lambert et al., 2023) and research attention to practices for creating and curating diverse, representative, and quality human feedback data (Feffer et al., 2023a; Kirk et al., 2024; Prabhakaran et al., 2021). However, less attention has been paid to human feedback collected through interactive features in interfaces to LLMs, in which users evaluate prompt completions of deployed systems. This form of feedback is captured after a given LLM has been trained and deployed, and such feedback may be used to improve future model iterations. For example, OpenAI used feedback on non-factual responses obtained through the feedback features in ChatGPT’s interface, with additional labelled comparison data, to train reward models that reduced GPT-4’s tendency to “hallucinate” compared to GPT-3.5 (OpenAI, 2023).
@@
The ChatGPT interface was also equipped with several interactive features designed to collect user feedback, allowing users to give thumbs-up/thumbs-down (see Fig. 1) and open-text feedback about responses (see Fig. 2 and Fig. 3), as well as to regenerate a response and identify whether the second response was better, the same, or worse than the original (see Fig. 4). Additionally, checkboxes in the open-text feedback windows allowed users to structure their feedback against the prominent “helpful, honest, and harmless” framework for AI alignment (Askell et al., 2021). Other interfaces to LLMs, such as Anthropic’s Claude.ai and Google’s Bard/Gemini, incorporate feedback features similar to the ChatGPT interface to gather user input and preferences. The deployment of LLMs through these interfaces has made these systems accessible to a vast and diverse user base, presenting new challenges related to scale.
Figure 1. Thumbs-up or down feedback - ChatGPT (May 2023)
[…]
An image of the ChatGPT interface showing a prompt completion. To the right of ChatGPT’s response, there are two buttons side by side: a thumbs-up icon on the left and a thumbs-down icon on the right. This represents the simple binary feedback mechanism offered to users after each prompt completion.
Figure 2. Open-text window after thumbs-up - ChatGPT (May 2023)
[…]
An image of a pop-up window in the ChatGPT interface that appears after a user gives a thumbs-up rating. The window is titled “Provide additional feedback” and contains a large text box for users to input detailed feedback, in which the following text appears: “What do you like about the response?”. Below the text box to the right is a “Submit feedback” button.
Figure 3. Open-text window after thumbs-down - ChatGPT (May 2023)
[…]
An image of a pop-up window in the ChatGPT interface that appears after a user gives a thumbs-down rating. The window is titled “Provide additional feedback” and contains a large text box for users to input detailed feedback, in which the following text appears: “What was the issue with the response? How could it be improved?”. Below the text box are three checkboxes labelled “This is harmful/unsafe”, “This isn’t true”, and “This isn’t helpful”. At the bottom right of the window is a “Submit feedback” button.
Figure 4. Feedback after regenerating a response - ChatGPT (May 2023)
[…]
A screenshot of the ChatGPT interface showing a regenerated response. Between the responses is a set of three buttons arranged horizontally, allowing the user to rate whether the new response is “Better”, “Worse”, or the “Same” compared to the original response.
@@ 2.2.1. Scale and Affordances
The standardisation of feedback features across interfaces to different systems exemplifies the proceduralisation and replication that characterise other digital systems deployed at large scale (Gillespie, 2020). LLMs are based on deep neural networks that learn from massive datasets, and these systems have grown significantly in size in recent years—for example, from millions of parameters in GPT-1 to billions in GPT-2 (Solaiman et al., 2019) and GPT-3 (Brown et al., 2020), and (reportedly) trillions in GPT-4 (Patel and Wong, 2023). But more than simply expanding the size of a system, scaling digital systems across global contexts invariably requires imposing a standardised worldview and organising phenomena that can operate context-independently (Tsing, 2012; Hanna and Park, 2020). This sameness is achieved through massive infrastructural investments, such as data centres, standardised software and policies, and global distribution networks (Young et al., 2024). Feedback features in interfaces to LLMs (and the associated infrastructure for managing, storing and processing feedback data) represent one such infrastructural investment designed to elicit participation at scale. Notably, these feedback features are identical for all users of a given interface to an LLM (e.g., ChatGPT, Bard/Gemini, or Claude.ai) and similar across different interfaces to various LLMs, despite the diversity of global contexts in which users interact with these systems.
However, this sameness can obscure the power dynamics inherent in the feedback process. Achieving scale in digital contexts is an aggregation of power—engineering the suppression of context to render complex social realities coherent through a computationally legible frame (Tsing, 2012; Gillespie, 2020). As Eubanks (Eubanks, 2017) demonstrates, the design of scaled technical systems can reinforce existing inequalities and marginalise certain voices. For example, people less likely to engage with these interfaces, such as those with limited digital literacy or access, may be underrepresented in the feedback data used to refine the family of models underlying the interfaces. The standardisation of feedback features across different systems might also lead to the homogenisation of user input, risking the amplification of certain biases or the suppression of diverse perspectives. It is thus imperative to critically examine what kinds of perspectives and contexts these feedback features make visible and valuable as forms of participation—both to those who participate and to those who integrate feedback from participants into ML workflows. The concept of affordances provides one useful lens for conducting such an analysis.
@@ 2.2.2. Mechanisms and Conditions Framework of Affordances
Affordance theory states that an animal perceives its environment not only in terms of the shapes of objects or the space between objects but also in the possibilities of objects for action (Gibson, 1977). Animals, to which Gibson (Gibson, 1977) refers, share the same environment but perceive it differently according to their affordances—their capacity to benefit or experience harm from a particular feature of the environment. Norman (Norman, 2008) later adapted affordance theory for technology design, arguing that technological objects afford different possibilities for humans and other agents and that those possibilities cannot be defined except by reference to the object’s adoption and use (Hildebrandt, 2015). The affordances of a technological object are thus the link between the object’s features and the outcomes of the technological object’s adoption and use by subjects (Davis, 2020). Davis and Chouinard (Davis and Chouinard, 2016) describe these links as “relational processes among users, designers, environments, and things”.
@@
Davis (Davis, 2020) operationalises the concept of affordances through a framework of mechanisms and conditions. Mechanisms describe the link between features of technological objects and the outcomes of their use by subjects, while conditions describe how affordances vary across different subjects and contexts. The framework proposes six mechanisms that contribute to the operationalisation of affordances: request, demand, encourage, discourage, refuse, and allow. These mechanisms are intended to serve as a tool for critical analysis rather than as a taxonomy of the features of technological objects, as each mechanism may reveal different aspects of a technical feature. Technological objects request and demand how user-subjects interact with them. Technological objects also encourage and discourage, refuse and allow actions of user-subjects. These mechanisms operate through three interrelated conditions: perception (how a subject perceives the artifact based on their skill and technical literacy), dexterity (what a subject can do with the artifact), and cultural and institutional legitimacy (the social and structural position of the user within institutions of power). The framework addresses binarism (assuming an artifact affords some function or not) and universalism (assuming an artifact operates identically for all) in affordance theory (Davis and Chouinard, 2016). Using this framework, we can explore “how, for whom, and under what circumstances” artefacts afford action (Davis, 2020).
@@
If we recognise ML-enabled technologies as material artefacts that reflect and affect users through design decisions, and acknowledge their potential to exacerbate inequality in social settings and workplaces (Davis, 2023), we must scrutinise the affordances of such technologies and applications. The mechanisms and conditions framework provides a structured approach to analyse how the design of feedback features in interfaces to LLMs affords certain actions while constraining others, thereby shaping the nature and scope of user input and participation in RLHF workflows. Affordance theory is useful for analysing how users generate evaluative feedback through LLM interfaces and how this feedback is subsequently shaped and constrained to function as cybernetic input in RLHF workflows. The framework’s emphasis on the interdependent and power-laden relationship between humans and technologies supports our critical examination of the sameness of scaled systems. Furthermore, the framework offers a vocabulary for the (re)design of feedback features to address issues we may identify for user input and public participation through socio-technical analysis (Davis, 2020).

### v2

@@ Abstract.
Large language models (LLMs) are now accessible to anyone with a computer, a web browser, and an internet connection via browser-based interfaces, shifting the dynamics of participation in AI development. This article examines how interactive feedback features in ChatGPT’s interface afford user participation in LLM iteration. Drawing on a survey of early ChatGPT users and applying the mechanisms and conditions framework of affordances, we analyse how these features shape user input. Our analysis indicates that these features encourage simple, frequent, and performance-focused feedback while discouraging collective input and discussions among users. Drawing on participatory design literature, we argue such constraints, if replicated across broader user bases, risk reinforcing power imbalances between users, the public, and companies developing LLMs. Our analysis contributes to the growing literature on participatory AI by critically examining the limitations of existing feedback processes and proposing directions for redesign. Rather than focusing solely on aligning model outputs with specific user preferences, we advocate for creating infrastructure that supports sustained dialogue about the purpose and applications of LLMs. This approach requires attention to the ongoing work of “infrastructuring”—creating and sustaining the social, technical, and institutional structures necessary to address matters of concern to stakeholders impacted by LLM development and deployment.
@@
journal: JRCjournalvolume: 2journalnumber: 4article: 16publicationmonth: 10journalyear: 2025doi: 10.1145/3748516ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Artificial intelligenceccs: Social and professional topics User characteristicsccs: Applied computing Sociology;
@@ 1. Introduction
The deployment of ChatGPT by OpenAI in 2022 marked a significant shift in how large language models (LLMs) are made available to the public. ChatGPT’s browser-based interface facilitates broad access to interactions with LLMs through text prompts, expanding user participation in the iterative development of LLMs. Previously, companies such as OpenAI and Google took a more cautious approach. OpenAI used application programming interfaces (APIs) and an iterative approach to deploy the GPT-2 and GPT-3 family of models (Solaiman et al., 2019; Brundage et al., 2022), limiting access to developers and researchers. Similarly, Google announced versions of the Language Model for Dialogue Applications (LaMDA) in 2021 and 2022 but did not make it publicly available through an interface or model release (Thoppilan et al., 2022). Following ChatGPT’s deployment, numerous other LLMs were made accessible via browser-based interfaces (e.g., Bard/Gemini and Claude.ai) or open releases (e.g., LLaMA).
@@
    What feedback features are present in the ChatGPT interface, and how are they used?
@@
    How do these feedback features, through specific mechanisms and conditions, afford particular types of user participation in the iteration of LLMs?
To address these questions, we first surveyed ChatGPT users who have submitted feedback. We then drew on the survey insights to analyse feedback features in the ChatGPT interface that allow users to submit feedback, using Davis’s (Davis, 2020) mechanisms and conditions framework of affordances. Understanding early adopter use of these feedback features offers insights into their initial affordances and potential limitations when scaling to broader populations; these early patterns are likely to shape the norms and standards for participation in LLM iteration. Our analysis indicates that the ChatGPT interface encourages simple, frequent, individualised, and unidirectional feedback on performance-related concerns while discouraging feedback from collectives or discussions among users, and refuses feedback modification over time. Further, technical background and engagement frequency significantly influence participation patterns.
We discuss implications arising from the constraints on user participation observed in our study for (re)designing feedback for LLMs. First, we explore deliberative approaches that might enable bidirectional communication between users and developers, changes in perspectives over time, and discussion about social or ethical issues among groups of users. Second, we consider the “infrastructuring” (Le Dantec and DiSalvo, 2013) that could support more expansive feedback channels, to enable sustained dialogue with stakeholders affected by LLM deployments about the purposes and applications of these technologies. Moving beyond simple preference elicitation toward more expansive feedback channels appears increasingly important as LLMs become ubiquitous and unquestionable infrastructure—systems embedded in other structures, transparent in use, and only visible upon breakdown (Star and Ruhleder, 1996).
@@ 2.1. Alignment and Participatory Approaches
Interest in participatory approaches for aligning LLMs is part of a broader participatory turn in AI research, design, and development (Delgado et al., 2023). Participatory AI projects involve end users and those impacted by AI systems in the design and development of those systems (Cooper and Zafiroglu, 2024), drawing on long-standing traditions in other fields, such as participatory design (Muller and Kuhn, 1993; Sundblad, 2011) and participatory planning (Arnstein, 1969; Kelty, 2020). Participatory AI projects now feature in articles, panels and workshops at prominent computer science conferences (Kulynych et al., 2020; Zytko et al., 2022) and in recent policy frameworks (Farthing et al., 2021; U.S. Office of Science and Technology Policy, 2022; Tabassi, 2023).
The participatory turn is one response to calls for moves away from formal models of fairness (Green, 2022) toward situated collaborations with communities affected by AI systems (Sloane, 2022). These calls reflect a growing recognition that algorithmic fairness is not merely a technical problem amenable to formal solutions, but a socio-political challenge requiring attention to structural inequities (Hoffmann, 2019; Wong, 2020). Indeed, technical approaches alone may obscure or reinforce structural inequities, leading some to advocate for alternatives to algorithmic fairness such as “algorithmic reparation” (Davis et al., 2021).
Participatory approaches promise more “democratic” AI systems (Kulynych et al., 2020) and more “pluralistic” alignment (Kirk et al., 2024). However, there is a broad spectrum of participatory practices in AI design, from informing or consulting participants to equitable partnerships with participants (Birhane et al., 2022; Corbett et al., 2023; Delgado et al., 2023). For the development of large-scale AI systems, participation has often taken passive and involuntary forms through the representation of groups of individuals in datasets used to train models (Denton et al., 2020; Sloane et al., 2022). While feedback features in interfaces to LLMs provide a more active and voluntary form of user participation, their emergence as the primary mode for publics to influence development processes has raised a question about the conceptual leap from structured preference elicitation to genuine participatory engagement (Feffer et al., 2023b).
Additionally, there is a significant gap between the scale of participatory methods in computing research and the scale at which recent AI systems are trained and deployed (Diaz and Madaio, 2025). Surveys of participatory approaches in computing research (Cooper et al., 2022) and participatory AI projects (Delgado et al., 2023) indicate that engagements have primarily been localised, focusing on one or a few groups or communities with shared characteristics. In contrast, recent deployments of LLMs through browser-based interfaces are accessible to anyone with access to a computer, a web browser, and an internet connection.
@@ 2.2. Interactive Feedback Features in Interfaces to LLMs
Feedback has a rich history rooted in cybernetics, which introduced the idea of feedback loops, where the output of a system is fed back as input, and queried the implications of feedback loops for modelling dynamical systems over time (Wiener, 1948). The concept of feedback loops influenced subsequent AI research directions (Minsky, 1961) and remains fundamental to AI systems that learn over time. Concurrently, “feedback” has also become a general term referring to information provided by an observer (e.g., users or other stakeholders) assessing a system’s performance.
Reinforcement Learning from Human Feedback (RLHF) combines these feedback types in modern AI development. RLHF begins with evaluative feedback, when a model generates two responses to a given prompt and human raters determine which response is better than the other according to some criteria. These human assessments are then used to create a cybernetic feedback loop: an RL algorithm uses the assessments to predict how a human would rate other responses, and an LLM is fine-tuned on the resulting reward model (Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022). While central to aligning pre-trained LLMs with specific preferences (Lambert et al., 2023), integrating human feedback into the learning process increases training time and labour costs (Maslej et al., 2024). RLHF requires large, diverse groups of human raters to assess prompt completions before fine-tuning the model (Prabhakaran et al., 2021).
The growing importance of RLHF has increased demand for human feedback data (Lambert et al., 2023) and research attention on practices for creating and curating diverse, representative, and quality human feedback data (Feffer et al., 2023a; Kirk et al., 2024; Prabhakaran et al., 2021). However, less attention has been paid to human feedback collected through interfaces to LLMs, in which users evaluate prompt completions of deployed systems. This post-deployment feedback may be used to improve future model iterations. For example, OpenAI used feedback on non-factual responses obtained through the feedback features in ChatGPT’s interface, with additional labelled comparison data, to train reward models that reduced GPT-4’s tendency to “hallucinate” compared to GPT-3.5 (OpenAI, 2023).
@@
Figure 1. Thumbs-up or down feedback - ChatGPT (May 2023).:
Figure 2. Open-text window after thumbs-up - ChatGPT (May 2023).:
The ChatGPT interface was also equipped with several interactive features, including: thumbs-up/thumbs-down ratings (see Figure 1) open-text feedback (see Figure 2 and Figure 3), and comparative feedback after a user regenerates a response (see Figure 4). Checkboxes in the open-text feedback windows allow users to structure their feedback against the prominent “helpful, honest, and harmless” framework for AI alignment (Askell et al., 2021). Similar features appear in other interfaces to LLMs, such as Google’s Bard/Gemini and Anthropic’s Claude.ai. The deployment of LLMs through these interfaces has made these systems accessible to a vast and diverse user base, presenting new challenges related to scale.
Figure 3. Open-text window after thumbs-down - ChatGPT (May 2023).:
Figure 4. Feedback after regenerating a response - ChatGPT (May 2023).:
@@ 2.2.1. Scale and Affordances
The standardisation of feedback features across interfaces to different systems exemplifies the proceduralisation and replication that characterise other digital systems deployed at large scale (Gillespie, 2020). LLMs have grown exponentially in size in recent years—from billions of parameters in GPT-2 (Solaiman et al., 2019) to (reportedly) trillions in GPT-4 (Patel and Wong, 2023). But more than simply expanding the size of a system, scaling digital systems across global contexts invariably requires imposing a standardised worldview and organising phenomena that can operate context-independently (Tsing, 2012; Hanna and Park, 2020). This sameness is achieved through massive infrastructural investments, such as data centres, standardised software and policies, and global distribution networks (Young et al., 2024). Feedback features (and the associated infrastructure for managing, storing, and processing feedback data) represent one such infrastructural investment designed to elicit participation at scale. Notably, these feedback features are identical for all users of a given interface to an LLM (e.g., ChatGPT, Bard/Gemini, or Claude.ai) and similar across different interfaces to various LLMs, despite the diversity of global contexts in which users interact with these systems.
However, this sameness can obscure the power dynamics inherent in the feedback process. Achieving scale in digital contexts is an aggregation of power—engineering the suppression of context to render complex social realities coherent through a computationally legible frame (Tsing, 2012; Gillespie, 2020). As Eubanks (Eubanks, 2017) demonstrates, the design of scaled technical systems can reinforce existing inequalities and marginalise certain voices. For example, people less likely to engage with these interfaces, such as those with limited digital literacy or access, may be underrepresented in the feedback data used to refine the family of models underlying the interfaces. The standardisation of feedback features across different systems might also lead to the homogenisation of user input, risking the amplification of certain biases or the suppression of diverse perspectives. The “helpful, honest, and harmless” framework (Askell et al., 2021), for example, is commonly used to structure feedback for RLHF workflows. However, applying such standardised frameworks risks imposing particular values across diverse contexts (Gabriel and Keeling, 2025). Definitions of “helpfulness” can vary culturally (Sambasivan et al., 2021), and “honesty” privileges certain epistemic standards inherent in a model’s training data (Denton et al., 2020), while determining “harmlessness” involves complex and contested judgments about potentialharms (Weidinger et al., 2022; Hoffmann, 2019; Gabriel and Keeling, 2025).
It is thus imperative to critically examine what kinds of perspectives and contexts these feedback features make visible and valuable as forms of participation—both to those who participate and to those who integrate feedback from participants into ML workflows. The concept of affordances provides one useful lens for conducting such an analysis.
@@ 2.2.2. Mechanisms and Conditions Framework of Affordances
Affordance theory states that an animal perceives its environment not only in terms of the shapes of objects or the space between objects but also in the possibilities of objects for action (Gibson, 1977). Animals share the same environment but perceive it differently according to their affordances—their capacity to benefit or experience harm from a particular feature of the environment. Norman (Norman, 2008) later adapted affordance theory for technology design, arguing that technological objects afford different possibilities for humans and other agents and that those possibilities cannot be defined except by reference to the object’s adoption and use (Hildebrandt, 2015). The affordances of a technological object are thus the link between the object’s features and the outcomes of the technological object’s adoption and use by subjects (Davis, 2020). Davis and Chouinard (2016) describe these links as “relational processes among users, designers, environments, and things”.
Davis (2020) operationalises the concept of affordances through a framework of mechanisms and conditions. The framework extends affordance theory by reframing the fundamental question away from what technologies afford to “how technologies afford, for whom, and under what circumstances(Davis, 2020). Mechanisms describe the link between features of technological objects and the outcomes of their use by subjects, while conditions describe how affordances vary across different subjects and contexts.
The framework proposes six mechanisms that contribute to the operationalisation of affordances: request, demand, encourage, discourage, refuse, and allow (Davis, 2020). These mechanisms are intended to serve as a tool for critical analysis rather than as a taxonomy of the features of technological objects, as each mechanism may reveal different aspects of a technical feature. Technological objects request and demand how user-subjects interact with them. Requests refer to bids placed by the artifact initiating some action or set of actions, while demands represent more forceful bids that leave users with little room for non-compliance. Technological objects also encourage and discourage, refuse and allow actions of user-subjects. These latter mechanisms represent responses from the artifact to activities that users seek to enact. Allow functions neutrally, applying to bids placed by both the artifact and user-subjects engaging with it. Importantly, these mechanisms are not mutually exclusive but operate simultaneously—features that request certain feedback formats may encourage some user behaviour at the same time as discouraging others.
@@
The mechanisms operate through three interrelated conditions: perception (how a subject perceives the artifact based on their skill and technical literacy), dexterity (what a subject can do with the artifact), and cultural and institutional legitimacy (the social and structural position of the user within institutions of power) (Davis, 2020). The interaction between mechanisms and conditions creates variations in how technologies afford actions for different users. For example, without sufficient dexterity or the requisite cultural and institutional legitimacy, what may be allowed becomes discouraged or refused (Davis, 2023). The framework thus addresses binarism (assuming an artifact affords some function or not) and universalism (assuming an artifact operates identically for all) in affordance theory (Davis and Chouinard, 2016).
@@
If we recognise ML-enabled technologies as material artefacts that reflect and affect users through design decisions, and acknowledge their potential to exacerbate inequality in social settings and workplaces (Davis, 2023), we must scrutinise the affordances of such technologies and applications. The mechanisms and conditions framework provides a structured approach to analyse how the design of feedback features in interfaces to LLMs affords certain actions while constraining others, thereby shaping the nature and scope of user input and participation in RLHF workflows. Furthermore, the framework offers a vocabulary for the (re)design of feedback features to address issues we may identify for user input and public participation through socio-technical analysis (Davis, 2020).